Determine the pH of a 0.268M solution of hydrocyanic acid. Express your answer to 2 decimals. Use the Ka value listed in your data booklet.
Answer #1

(ka) HON = 6.2x 10-10 given: 0.268 m Solution HON H+ + CN- ka= [at] [CNO [HCM 6. 2x10-10 = (2) (a) => 32+ 6.2X100% - 1.662x10go Now Pok= - pag [on] z - Toy 1.23x105] - 4.92 pH = 14-4.92 9H = 9.1 Dus Scanned with CamScanner